OpenStudy (anonymous):

Find all solutions of the equation in the interval [0,2pi) algebraically. Use the table feature of a graphing utility to check your answers numerically. sin^2x + cos x + 1=0

OpenStudy (anonymous):

Does the identity sin^2(x) =1-cos^2(x) help you with anything

OpenStudy (anonymous):

no haha

OpenStudy (jdoe0001):

\(\bf sin^2(x) + cos(x) + 1=0\\ \textrm{as Ahmad1 said, using }sin^2(x) = 1-cos^2(x)\\ 1-cos^2(x)+ cos(x) + 1=0 \implies -cos^2(x)+ cos(x) +2 = 0\)

OpenStudy (jdoe0001):

so you have a quadratic equation

OpenStudy (anonymous):

and now you can substitute cos(x)=u and solve for u , and substitute back after that to, get the answer in terms of cosine , but just make sure that x stays in the interval you stated..

OpenStudy (anonymous):

ummmmm can someone just write the answer lol

OpenStudy (anonymous):

(cos(x)-2)(cos(x)+1)=0 either cos(x)=2 and that's not possible, or cos(x)=-1 then x=pi

OpenStudy (anonymous):

I just factored the expression ..